OpenStudy (anonymous):

t(1)=2 t(2)=3 t(3)=t(1)*t(2)+1 t(4)=t(1)*t(2)*t(3)+1 . . . t(n)=t(1)*t(2)*t(3)*...*t(n-1)+1 PROVE or DISPROVE that t(n) will surely be PRIME

OpenStudy (turingtest):

This sounds like a very hard problem.

OpenStudy (anonymous):

Maybe not....

OpenStudy (anonymous):

I dont know..... But I think it will not be surely prime.

OpenStudy (anonymous):

If not, we just need a counterexample.

OpenStudy (turingtest):

I doubt it is all primes as well.

OpenStudy (unklerhaukus):

remainder is gonna be one

OpenStudy (klimenkov):

Looks like an Euclid proof of the infinite number of prime numbers.

OpenStudy (anonymous):

Euclid never said there was an infinity of primes (didn't believe in infinity)

OpenStudy (swissgirl):

t(5)=2*3*7*43+1=1807 1807/13=139

OpenStudy (anonymous):

Euclid said that you could always construct another one out of a supposedly complete list.

OpenStudy (unklerhaukus):

that dosent make sense

OpenStudy (klimenkov):

There are infinitely many primes, as demonstrated by Euclid around 300 BC. http://en.wikipedia.org/wiki/Prime_number

OpenStudy (anonymous):

True but he never said anything about infinity (Greeks weren't too keen on that idea)

OpenStudy (klimenkov):

You deepened into the history. But I spoke about the method.

OpenStudy (anonymous):

The method, I agree, is very like the question.....

OpenStudy (anonymous):

P_n = p1p2p3....+1

OpenStudy (anonymous):

The "infinitely many" part got added later.....

OpenStudy (anonymous):

Personally, I like "you can always get another one" better....

OpenStudy (anonymous):

"Construct another one"

OpenStudy (unklerhaukus):

do you mean induction

OpenStudy (anonymous):

No, it is an explicit construction...

OpenStudy (anonymous):

You give me a list of primes and say "That's all there are" And I give you another one not in the list...